If a= 2x-1 and b= 3-x, which of the following expresses a in term of b?
a. 5-2b
b. 6-b
c. b-5
d. 2b-6
e. 2b-5

Answer by stanbon(75887)

If a= 2x-1 and b= 3-x, which of the following expresses a in term of b?
-----
a = 2x-1 = -2(3-x)+5 = -2b+5
-----
Cheers,
Stan H.
------------
a. 5-2b
b. 6-b
c. b-5
d. 2b-6
e. 2b-5